factorize 8x^3+5x^3+1

factorize

8x^3+5x^3+1

Grade:10

1 Answers

Sudheesh Singanamalla
13 years ago

8x^3+5x^3+1?

13 x^3 + 1

using a3 + b3  = (a + b)(a2 – ab + b2)

a=(13)1/3x , b=1

so

13x^3 + 1 = ((13)1/3x+1)((13)2/3 - (13)1/3x + 1)

 

Now find the roots

 roots are

x = - 1 / (13)1/3

x = (13)1/3+√(13)1/2- 52*(13)1/3

x = (13)1/3+√(13)1/2+ 52*(13)1/3

----------------------------------------------------------------

PLEASE APPROVE

----------------------------------------------------------------